OpenStudy (anonymous):

sqrt 9 m^2 n^2 + 2 sqrt m^2 n^2 - 3mn sqrt mn 2mn

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@24/7

Nnesha (nnesha):

exponent rules! \[\huge\rm \sqrt[n]{x^m} = x^\frac{ m }{ n }\] change root to exponent form